Ticket Number: HEA-NOC/20070206-10 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: EPA
Ticket Opened: 20070206 14:32 UTC Problem Start: 20070206 14:26 UTC
Ticket Closed: 20070206 16:13 UTC Problem End: 20070206 16:12 UTC
Notification was received that the firewall in the EPA is to be rebooted
Connectivity to the EPA from HEAnet.
20070206 14:26 BS took call from Frank O'Brien. Opening ticket to track.
20070206 16:13 AH EPA confirm work was completed successfully.
This ticket can be monitored at http://www.hea.net/tickets/20070206-10